OpenStudy (anonymous):

the uniform acceleration with which a man could start running from rest behind a bus that goes past him with a constant acceleration of 5m/s on a straight road to catch it in 20m distance from start is ?

OpenStudy (anonymous):

the ans=2.5m/s^2

OpenStudy (ajprincess):

the bus travels with constant acceleration ir constant velocity?

OpenStudy (ajprincess):

*or

OpenStudy (anonymous):

constant speed

OpenStudy (anonymous):

a constant acceleration this is wrong it is contant speed

OpenStudy (anonymous):

sorry for the mistake

OpenStudy (ajprincess):

Since the bus travels with uniform speed use s=ut 20=5t t=4 use s=ut+1/2at^2 for the man here t will be 4s initial velocity is 0 bcoz the person starts frm rest. 20=1/2*a*16 20=8a a=2.5

OpenStudy (anonymous):

thanzz